The probability that a first-time tourist to the city of
Toledo will visit the Art Museum is 0.5, will visit the Toledo
Zoo is 0.4, and will visit both is 0.25. If a first-time tourist
to Toledo is randomly selected, find the probability that the
tourist will visit:
(a) the Art Museum or the Toledo Zoo
(b) neither of these attractions, and
(c) one, but not both, of these attractions.

Answer by math_tutor2020(3816)


M = visit museum
Z = visit zoo
P(M) = 0.5
P(Z) = 0.4
P(M and Z) = 0.25

Part (a)
P(M or Z) = P(M) + P(Z) - P(M and Z) ........ inclusion-exclusion principle
= 0.5+0.4-0.25
= 0.65

Part (b)
P(neither M nor Z) = 1 - P(M or Z)
= 1 - 0.65
= 0.35

Part (c)
P(M or Z but not both) = P(M or Z) - P(M and Z)
= 0.65-0.25
= 0.4

--------------------------------------------------------------------------

Summary of answers
(a) 0.65
(b) 0.35
(c) 0.4